A 37.98 g solid sample of mercury is heated from -39 oC its melting point, completely melted then heated to 106 oC, How much energy was absorbed by the sample? (Answer in kJ)

S=0.14 J/goC Delta Hf = 0.011 kJ/g



1
Expert's answer
2021-03-29T06:04:30-0400

Solution:

This problem can be summarized thusly:

1) A sample of solid mercury is melted at -39°C. The heat absorbed is calculated by multiplying the grams of solid mercury by the heat of fusion (ΔHf) in kJ/g.

q1 = ΔHf × m = 0.011 kJ/g × 37.98 g = 0.418 kJ

q1 = 0.418 kJ


2) A sample of liquid mercury is heated from -39°C to 106°C. The heat absorbed is calculated by using the specific heat of liquid mercury (S = 0.14 J/g°C) and the equation: q = m × S × ΔT.

q2 = m × S × (Tf - Ti)

q2 = 37.98 g × 0.14 J/g°C × (106 - (-39))°C = 770.99 J = 771 J = 0.771 kJ

q2 = 0.771 kJ


q = q1 + q2 = 0.418 kJ + 0.771 kJ = 1.189 kJ = 1.19 kJ

q = 1.19 kJ


Answer: 1.19 kJ of energy was absorbed by the sample.